OpenStudy (anonymous):

find the difference quotient of f; that is, find f(x+h)-f(x) over h. f(x)= 4x+3

OpenStudy (anonymous):

wrong problem. its f(x)= x^2-x+4

hero (hero):

Oh cool, I get to do all of that all over again....

OpenStudy (anonymous):

the answer is f ' (x) f(x) = x^2 - x + 4 f ' (x) = 2x - 1

hero (hero):

NO, difference quotient is not the same as derivative

hero (hero):

The limit of the difference quotient as h goes to zero is the derivative....but not the difference quotient itself

hero (hero):

I don't feel like re-writing all of that all over again

OpenStudy (anonymous):

well get off.. i need somebody that can help m

OpenStudy (anonymous):

me*

hero (hero):

Get off? Am I on a train? lol

hero (hero):

Is that how you normally talk to people attempting to help you?